OpenStudy (anonymous):

How do you write this equation? The sum of three numbers is eighty-six. The first of the three numbers is four less than two times the second. The third number is six more than the first. Find the numbers. The correct 3 numbers are x=31.2; y=17.6; z=37.2 How do you get these?

OpenStudy (flvs.net):

What do you think?

OpenStudy (anonymous):

Well I thought, x+y+z=86, then I would uses y as my reference point. 2y-2+y+6(2y-2)=86 BUT this is so wrong. Please help!

OpenStudy (welshfella):

let x be the second number then first number = 2x - 4 third number = 2x - 4 + 6 = 2x + 2 the sum is 86 therefore we have x + 2x -4 + 2x + 2 = 86 solve for x